A contractor mixes concrete from bags of pre-mix. How many bags with 10% cement should he mix with 8 bags of 12.5 cement to produce a mix of 11% cement?
Please help - I'm stuck.

Answer by ptaylor(2198)

Let x=the number of bags of 10% cement
We know that the amt of pure cement in the unknown number of bags (x)(.10) plus amt of pure cement in the 8 bags (8)(.125) must equal the amt of pure cement in the final mixture (x+8)(.11). Thus, the equation to solve is:
.10x+(8)(.125)=(x+8)(.11) simplifying, we get:
.10x+1=.11x+.88 further simplifying, we have:
.01x=.12 or
x=12 bags
Hope this helps -----ptaylor